find the balance after 15 years of a $2500 investment earning 5.5% interest compouunded continuously
Answer by checkley77(12844)

I=2,500e^rt
I=2,500*2.71828^.055*15
I=2,500*2.71828^.825
I=2,500*2.282
I=$5,704.70 ans.